# Multi-PA: A Multi-perspective Benchmark on Privacy Assessment for Large Vision-Language Models

### v3

@@ Abstract
Large Vision-Language Models (LVLMs) exhibit impressive potential across various tasks but also face significant privacy risks, limiting their practical applications. Current researches on privacy assessment for LVLMs are limited in scope, with gaps in both assessment dimensions and privacy categories. To bridge this gap, we propose Multi-PA, a comprehensive benchmark for evaluating the privacy preservation capabilities of LVLMs in terms of privacy awareness and leakage. Privacy awareness measures the model’s ability to recognize the privacy sensitivity of input data, while privacy leakage assesses the risk of the model unintentionally disclosing privacy information in its output. We design a range of sub-tasks to thoroughly evaluate the model’s privacy protection offered by LVLMs. Multi-PA covers 26 categories of personal privacy, 15 categories of trade secrets, and 18 categories of state secrets, totaling 31,962 samples. Based on Multi-PA, we evaluate the privacy preservation capabilities of 21 open-source and 2 closed-source LVLMs. Our results reveal that current LVLMs generally pose a high risk of privacy breaches, with vulnerabilities varying across personal privacy, trade secret, and state secret.
@@ I Introduction
Since the emergence of ChatGPT [OpenAI], Large Language Models (LLMs) have attracted significant attention and become pivotal to the advancement of artificial intelligence. With the scaling up of training data and model parameters, the comprehension and reasoning capabilities of LLMs have improved substantially [achiam2023gpt, anil2023palm, jiang2023mistral, touvron2023llama]. Building upon these advancements, recent research has incorporated visual modalities into LLMs [bai2023qwen, chen2023minigpt, liu2024visual, team2023gemini, zhu2023minigpt, wang2023cogvlm], giving rise to Large Vision-Language Models (LVLMs). LVLMs typically consist of three essential components [yin2023survey]: an image encoder, a text encoder, and a strategy for aligning the information from both encoders. LVLMs demonstrate outstanding performance across a variety of tasks, including image captioning, visual question answering, etc.
@@
Despite their impressive performance in various tasks, the privacy risks associated with LVLMs remain a significant concern [principles2017future]. The improper use or leakage of private data in training set [nasr2023scalable] may lead to severe privacy violations, raising both legal and ethical concerns. In recent years, data protection regulations have become increasingly stringent. For example, laws like GDPR [gdprinfoGeneralData] require the developers to address privacy issues throughout the data lifecycle. In this scenario, researchers have begun to adopt techniques such as differential privacy[dwork2006differential, 10795202, 10806731, abadi2016deep] and federated learning[mammen2021federated, 10731856, 10741585, 10793087, chen2024integration] to reduce the privacy risks of models. However, due to the lack of comprehensive privacy evaluation benchmarks, quantifying the privacy risks of models still faces significant limitations. Comprehensive privacy assessments are essential to identify potential risks during both the model training and inference, facilitating the design of the protective measures. Although previous security assessments have considered privacy risks [zhang2024benchmarking, gu2025mllmguard, wang2024cross, sun2024trustllm], they primarily concentrate on personal privacy, overlooking other sensitive categories such as trade secret and state secret. Additionally, these assessments generally focus on quantifying the extent of memory data leakage from the model, neglecting the evaluation of the model’s ability to perceive and infer privacy-related information from the users’ inputs [tomekcce2025private]. As shown in Figure 1, different fundamental capabilities of VLMs correspond to distinct perspectives on privacy leakage. Analyzing these perspectives separately allows for a clearer identification of the model’s potential privacy risks.
@@
To address the aforementioned challenges, we establish Multi-PA, a comprehensive benchmark for evaluating privacy risks in Large Vision-Language Models (LVLMs). Inspired by TrustLLM [sun2024trustllm] and MultiTrust [zhang2024benchmarking], we assess the privacy risks along two dimensions: Privacy Awareness and Privacy Leakage, as illustrated in Figure 2. The former measures the model’s ability to recognize risks associated with privacy violations, while the latter evaluates the extent to which the model unintentionally discloses privacy-related information. For Privacy Awareness, we design three specific tasks: Privacy Image Recognition, Privacy Question Detection, and Privacy InfoFlow Assessment. For Privacy Leakage, we categorize tasks based on the model’s capacities: 1) Perception Leakage (extracting privacy-related information), 2) Reasoning Leakage (inferring privacy-related information), 3) Memory Leakage (disclosing privacy-related information from training data). Multi-PA encompasses 59 categories of privacy, offering a comprehensive evaluation framework for privacy risks. Each privacy category is associated with carefully defined privacy/non-privacy attributes, which serve as question-asking targets. These attributes are paired with relevant images to create Visual Question Answering (VQA) samples, yielding a total of 31,962 samples. Based on Multi-PA, we conduct extensive privacy risk assessments on 21 open-source and 2 closed-source models. To avoid excessively conservative models from achieving improper ranks, we propose Expect-to-Answer ($EtA$), which balances the model’s tendency to refuse responses to privacy-sensitive questions with its responsiveness to questions involving non-sensitive attributes. Our findings reveal that current LVLMs generally pose a significant risk of privacy leakage and exhibit varying degrees of vulnerability across different privacy types. Multi-PA offers a vital benchmark for assessing the privacy risks of LVLMs, encouraging further research on enhancing their privacy safeguards. Our contributions are as follows:
@@
  • We evaluate 21 open-source models and 2 closed-source models on Multi-PA, uncovering significant privacy risks in current LVLMs. These findings provide new insights for advancing the development of privacy enhancement.
@@ II Related work
Benchmark Privacy Awareness Privacy Leakage Personal Privacy Trade Secret State Secret Models Samples Sensitive Insensitive CONFAIDE [mireshghallah2023can] $\times$ $\times$ 10 N/A N/A 6 766 VIP dataset [tomekcce2025private] $\times$ $\times$ 8 N/A N/A 7 554 MLLMGUARD [gu2025mllmguard] $\times$ $\times$ - - - 13 323 MultiTrust [zhang2024benchmarking] $\times$ 16 N/A N/A 21 3,415 Multi-$\text{P}^{\text{2}}$A (Ours) 26 15 18 23 31,962
@@ II-A Large Vision-Language Models
Large Vision-Language Models (LVLMs) represent a significant advancement at the intersection of natural language processing and computer vision, allowing for the simultaneous handling of both visual and textual inputs. LVLMs, such as LLaVA-1.5 [liu2024visual], BLIP2 [li2023blip], MiniGPT-4[zhu2023minigpt], MiniGPT-v2 [chen2023minigpt], Otter [li2023ottermultimodalmodelincontext], Instruct-BLIP [dai2023instructblipgeneralpurposevisionlanguagemodels], and InternLM-XComposer [internlmxcomposer], facilitate visual-text interactions that support complex multimodal reasoning, thereby enhancing performance across diverse tasks. Recently, new models, including MiniCPM-LLaMA-v2.5 [yao2024minicpm], mPLUG-OWL2 [ye2023mplugowl2], Phi-3-Vision [abdin2024phi], Qwen-VL-Chat [Qwen-VL], Shikra [chen2023shikra], and Yi-VL [ai2024yi], introduce unique features and targeted applications, further diversifying LVLM capabilities. Notably, closed-source models like Gemini [team2023gemini] and GPT-4V [OpenAI-gpt4V] achieve expert-level performance in specific domains, underscoring the specialized strengths and practical utility of current LVLMs across various sectors.
@@ II-B Privacy Evaluation of Language Models
Currently, privacy evaluation benchmarks can be divided into instance-level benchmarks and category-level benchmarks. Instance-level benchmarks originate from the discovery that models can memorize training data and disclose such data during inference [carlini2022quantifying, carlini2021extracting, jayaraman2022active, yu2023bag, staab2023beyond]. Instance-level benchmarks, such as LLM-PBE [li2024llm] and P-Bench [li2023p], define privacy as specific samples within training data. They measure models’ privacy risks through the attack success rate (ASR) of privacy attacks, like data extraction attack (DEA) [nasr2023scalable, carlini2019secret, carlini2021extracting, mireshghallah2022empirical] and membership inference attack (MIA) [shokri2017membership, mattern2023membership, carlini2021extracting]. However, these benchmarks typically require researchers to access models’ training data and internal representations, which consequently restricts the applicability of these methods in various practical scenarios.
Instance-level benchmarks focus on recapitulation risks of training data, while category-level benchmarks evaluate models’ risks on outputting sensitive data. Specifically, category-level benchmarks evaluate the handling of models when processing privacy-related requests, where any compliance response to such requests constitutes a privacy violation, irrespective of the correctness of their output. To evaluate models’ privacy risks, category-level benchmarks may design a set of privacy-related questions for each predefined privacy category. Currently, visual question answering (VQA) formats are widely used for category-level privacy evaluation [shi2024assessment, mireshghallah2023can, staab2023beyond, tomekcce2025private, gu2025mllmguard, zhang2024benchmarking, sun2024trustllm]. CONFAIDE [mireshghallah2023can] employs psychological theories to evaluate model’s alignment with human judgment when evaluating the flow of privacy-related data. Some researches [staab2023beyond, tomekcce2025private] demonstrate that models can infer personal attributes from inputs, thereby infringing on personal privacy, and construct a benchmark to quantify the privacy reasoning capabilities of models. To further refine the scope of privacy protection, MLLMGuard [gu2025mllmguard] categorizes privacy types into personal privacy, trade secret, and state secret, but the limited dataset scale restricts the reliable of the evaluation. MultiTrust [zhang2024benchmarking] and TrustLLM [sun2024trustllm] assess privacy security of models from two dimensions: privacy awareness and privacy leakage, which improve the privacy assessment framework. As shown in Table I, we further expand the privacy scopes and dataset scale, constructing a more comprehensive evaluation benchmark. Furthermore, we restructure the evaluation framework with awareness and leakage by categorizing privacy leakage into granular subdimensions aligned with LVLM’s core competencies (perception, reasoning, and memorization). Finally, we design Insensitive Questions to evaluate the unintended impact of privacy protection mechanisms on benign query responsiveness.
@@ III-A Definition of Privacy
Past category-level privacy benchmarks lack a formalized privacy definition. Some category-level privacy benchmarks [mireshghallah2023can, tomekcce2025private, zhang2024benchmarking] focus on the security of personal identifiable information, which is only a subset of privacy. MLLMGUARD [gu2025mllmguard] expands the scope of privacy to personal privacy, trade and state secrets, but it still lacks a clear definition of privacy.
@@ IV Multi-PA Benchmark
Figure 3: VQA Generation Process. For security reasons, we obscure the private parts in images. We build Image Database and Attribute Database by collecting images and designing attributes for each privacy category. For each task, we create a variety of question templates which will be randomly selected to generate samples. Each VQA sample is the combination of an image from Image Database and a question from Question Templates. For Privacy Question Detection and Privacy InfoFlow Assessment, context of each sample is respectively from corresponding question in Privacy Leakage and sample in CONFAIDE [mireshghallah2023can].
@@ IV-A Dataset Overview
As shown in Table 2, we evaluate the privacy security of LVLMs from Privacy Awareness and Privacy Leakage.
@@
Privacy Question Detection. Privacy Question Detection assesses the model’s capacity to discern the privacy risks of input queries. Notably, an image with privacy-related visual cues may be paired with a benign question, while a privacy-unrelated image may be associated with a sensitive question (e.g., querying for a politician’s home address given their publicly available photo is undeniably inappropriate). Therefore, the model’s assessment on the privacy sensitive of input queries may be context-dependent. To comprehensively evaluate a model’s awareness on the privacy sensitivity of input questions, we construct a dataset comprising image-question pairs, where images may contain private visual cues or not, and questions may be either privacy-related or privacy-unrelated. The model is tasked with determining the privacy relevance of each query.
@@
Privacy Leakage. Privacy leakage occurs when a model, during its interaction with users, either intentionally or unintentionally reveals privacy-related data. We categorize privacy leakage into two paradigms: assistant-like privacy leakage and database-like privacy leakage. Assistant-like leakage arises from insecure requests, where the model is compelled to disclose private data based on the provided input images. Database-like leakage on the other hand, occurs when private information, potentially present in the model’s training data, is memorized later exposed during interactions. Furthermore, we divide assistant-like leakage into “Perception” Leakage and “Reasoning” Leakage based on the capacities the model uses in the leakage process, and using “Memory” Leakage to evaluate the model’s database-like leakage. Additionally, we generate a variety of requests that do not involve private information, recorded as “Insensitive Questions”, which are used to assess the potential negative impact of privacy protection mechanisms on the model’s responsiveness to benign inquiries.
@@
Reasoning Leakage. “Reasoning” Leakage evaluates the risk of LVLMs inferring privacy based on visual cues within input images. Recent researches have demonstrated language models can infer privacy-sensitive data from indirect clues [staab2023beyond, tomekcce2025private]. Motivated by these findings, we consider the ability of models to infer privacy attributes as a crucial aspect of model’s privacy risks. To prevent models from facilitating the illicit prediction of private data, a privacy-preserving model should avoid generalizing its reasoning abilities to inferring privacy attributes. Therefore, for each privacy category, we curate a series of inferable privacy attributes for each privacy category. These attributes can be indirectly deduced from cues present in input images. For example, in an image containing a company’s revenue data, the company’s financial status might be considered as an inferable attribute.
Memory Leakage. “Memory” Leakage evaluates the risk of LVLMs revealing privacy retained from its training process. Researches have shown that models can memorize training data and unintentionally disclose it during interactions [carlini2022quantifying, carlini2021extracting, jayaraman2022active, yu2023bag, staab2023beyond]. Privacy with long-term sensitivity, like PII of public figures, may be unintentionally exposed on social platforms and subsequently incorporated into model training datasets. The privacy leakage of such information may lead to detrimental consequences. To assess the risk of a model leaking private information from its training dataset, we define weakly associated privacy attributes for each privacy category. Weakly associated privacy attributes refer to information that cannot be directly or indirectly derived from the input data. However, models may still output these attributes based on private information it has memorized from its training dataset. For example, given images of some businessmen, we ask models about their home addresses. Such private data may have appeared in journalists’ reports. When these reports are used for model training, models might memorize such private data and provide correct answer during inference.

### v4

# arXiv:2412.19496v4[cs.CR] 02 Mar 2026

@@ Abstract
Large Vision-Language Models (LVLMs) exhibit impressive potential across various tasks but also face significant privacy risks, limiting their practical applications. Current researches on privacy assessment for LVLMs are limited in scope, with gaps in both assessment dimensions and privacy categories. To bridge this gap, we propose Multi-PA, a comprehensive benchmark for evaluating the privacy preservation capabilities of LVLMs in terms of privacy awareness and leakage. Privacy awareness measures the model’s ability to recognize the privacy sensitivity of input data, while privacy leakage assesses the risk of the model unintentionally disclosing privacy information in its output. We design a range of sub-tasks to thoroughly evaluate the model’s privacy protection offered by LVLMs. Multi-PA covers 26 categories of personal privacy, 15 categories of trade secrets, and 18 categories of state secrets, totaling 31,962 samples. Based on Multi-PA, we evaluate the privacy preservation capabilities of 21 open-source and 2 closed-source LVLMs. Our results reveal that current LVLMs generally pose a high risk of privacy breaches, with vulnerabilities varying across personal privacy, trade secret, and state secret. We release our code and data at: https://github.com/Xiangkui-Cao/Multi-P2A.
@@ I Introduction
Since the emergence of ChatGPT [31], Large Language Models (LLMs) have attracted significant attention and become pivotal to the advancement of artificial intelligence. Recent research has incorporated visual modalities into LLMs [3, 11, 23, 41, 56, 45], giving rise to Large Vision-Language Models (LVLMs). LVLMs typically consist of three essential components [51]: an image encoder, a text encoder, and a strategy for aligning the information from both encoders. LVLMs demonstrate outstanding performance across a variety of tasks, including image captioning, visual question answering, etc.
@@
Despite their impressive performance in various tasks, the privacy risks associated with LVLMs remain a significant concern [35]. The improper use or leakage of private data in training set [30] may lead to severe privacy violations, raising both legal and ethical concerns. As shown in Figure 1, different fundamental capabilities of VLMs correspond to distinct perspectives on privacy leakage. As these capabilities become increasingly comprehensive, privacy risks may arise from capability overreach along multiple dimensions, underscoring the need for holistic safeguards. In recent years, data protection regulations have become increasingly stringent. Laws like GDPR [15] require the developers to address privacy issues throughout the data lifecycle. In this scenario, researchers have begun to adopt techniques such as differential privacy[14, 44, 48, 1] and federated learning[26, 47, 46, 24, 10] to reduce the privacy risks of models. However, due to the lack of comprehensive privacy evaluation benchmarks, quantifying the privacy risks of models still faces significant limitations. VIP dataset [39] demonstrates that models can infer personal attributes from inputs and proposes a benchmark for evaluating privacy-related reasoning capabilities. However, it abstracts away privacy risks associated with perception and memory, which are also critical sources of privacy leakage in multimodal models. MultiTrust [55] and TrustLLM [40] divide privacy evaluation into two dimensions, namely privacy awareness and privacy leakage, thereby enriching the privacy assessment framework. However, these dimensions are further organized by task categories rather than by underlying model capabilities, which limits comprehensive coverage of privacy risks across all model abilities. In addition, privacy awareness and privacy leakage are treated independently, leaving their interaction largely unexplored. While existing benchmarks largely concentrate on personal privacy, MLLMGuard [17] expands privacy types to encompass trade secrets and state secrets. Nevertheless, the constrained dataset size hinders robust, type-wise analysis of privacy risks across models.
Motivated by these limitations, we seek to systematically characterize the privacy risks of current LVLMs from a holistic perspective. Rather than isolating individual aspects, our goal is to jointly examine how models perceive privacy, reason about sensitive information, and potentially disclose private content during realistic interactions. In particular, we focus on the alignment between privacy awareness and privacy leakage, the role of different model capabilities in shaping privacy risks, and the uneven protection across privacy types. Based on these considerations, we formulate the following research points:
  • P1 (Awareness-Leakage Discrepancy): The relationship between privacy awareness and privacy leakage in LVLMs remains insufficiently understood.
  • P2 (Capability-Specific Bias): Privacy risks may differ across a model’s capabilities, such as perception, reasoning, and memory. We investigate which capabilities are more prone to privacy failures and analyze the different sources of privacy leakage.
  • P3 (Privacy-Type Bias): Whether LVLMs prioritize certain privacy types over others, leading to uneven privacy protection across categories.
  • P4 (Defense Impact): In addition to affecting privacy leakage, privacy-enhanced strategies may alter models’ privacy awareness. How such strategies influence privacy awareness, and how these changes relate to privacy leakage, remains insufficiently understood.
@@
To systematically investigate these research points, we establish Multi-PA, a comprehensive benchmark for evaluating privacy risks in Large Vision-Language Models (LVLMs). Inspired by TrustLLM [40] and MultiTrust [55], we assess the privacy risks along two dimensions: Privacy Awareness and Privacy Leakage, as illustrated in Figure 2. The former measures the model’s ability to recognize risks associated with privacy violations, while the latter evaluates the extent to which the model unintentionally discloses privacy-related information. For Privacy Awareness, we design three specific tasks: Privacy Image Recognition, Privacy Question Detection, and Privacy InfoFlow Assessment. For Privacy Leakage, we categorize tasks based on the model’s capacities: 1) Perception Leakage (extracting privacy-related information), 2) Reasoning Leakage (inferring privacy-related information), 3) Memory Leakage (disclosing privacy-related information from training data). Multi-PA encompasses 59 categories of privacy, offering a comprehensive evaluation framework for privacy risks. Each privacy category is associated with carefully defined privacy/non-privacy attributes, which serve as question-asking targets. These attributes are paired with relevant images to create Visual Question Answering (VQA) samples, yielding a total of 31,962 samples. To avoid excessively conservative models from achieving improper ranks, we propose Expect-to-Answer ($EtA$), which balances the model’s tendency to refuse responses to privacy-sensitive questions with its responsiveness to questions involving non-sensitive attributes. Based on Multi-PA, we conduct extensive privacy risk assessments on 25 open-source and 2 closed-source models. Our contributions are as follows:
@@
  • We evaluate 25 open-source models and 2 closed-source models on Multi-PA, uncovering significant privacy risks in current LVLMs. These findings provide new insights for advancing the development of privacy enhancement.
@@ II Related work
Benchmark Privacy Awareness Privacy Leakage Personal Privacy Trade Secret State Secret Models Samples Sensitive Insensitive CONFAIDE [29] $\times$ $\times$ 10 N/A N/A 6 766 VIP dataset [43] $\times$ $\times$ 8 N/A N/A 7 554 MLLMGUARD [17] $\times$ $\times$ - - - 13 323 MultiTrust [55] $\times$ 16 N/A N/A 21 3,415 Multi-PA (Ours) 26 15 18 27 31,962
@@ II-A Large Vision-Language Models
Large Vision-Language Models (LVLMs) represent a significant advancement at the intersection of natural language processing and computer vision, allowing for the simultaneous handling of both visual and textual inputs. facilitate visual-text interactions that support complex multimodal reasoning, thereby enhancing performance across diverse tasks. Recent models, like Phi-3-Vision [2], Qwen-VL-Chat [3], introduce unique features and targeted applications, further diversifying LVLM capabilities. Notably, closed-source models like Gemini [41] and GPT-4V [32] achieve expert-level performance in specific domains, underscoring the specialized strengths and practical utility of current LVLMs across various sectors.
@@ II-B Privacy Evaluation of Language Models
Currently, privacy evaluation benchmarks can be divided into instance-level benchmarks and category-level benchmarks. Instance-level benchmarks originate from the discovery that models can memorize training data and disclose such data during inference [7, 9, 18, 53, 39]. Instance-level benchmarks, such as LLM-PBE [22] and P-Bench [20], define privacy as specific samples within training data. They measure models’ privacy risks through the attack success rate (ASR) of privacy attacks, like data extraction attack (DEA) [30, 8, 9, 28] and membership inference attack (MIA) [38, 27, 9]. However, these benchmarks typically require researchers to access models’ training data and internal representations, which consequently restricts the applicability of these methods in various practical scenarios.
In contrast, category-level benchmarks evaluate whether models produce sensitive outputs when responding to privacy-related requests. Under this formulation, any compliant response to a privacy-sensitive query is regarded as a privacy violation, regardless of output correctness. To operationalize this definition, category-level benchmarks typically construct privacy-related question sets for predefined privacy categories. Visual question answering (VQA) has become the dominant evaluation format for category-level privacy benchmarks, particularly in multimodal settings [37, 29, 39, 43, 17, 55, 40]. Within this line of work, CONFAIDE [29] incorporates psychological theories to assess the alignment between model outputs and human judgments regarding privacy-related information flow. Other studies show that models can infer personal attributes from inputs, thereby violating personal privacy, and propose benchmarks to evaluate such privacy reasoning behaviors [43].
Most existing category-level benchmarks primarily focus on personal privacy. MLLMGuard [17] extends the privacy scope to include trade secrets and state secrets, although its relatively limited dataset size constrains fine-grained analysis across privacy types. MultiTrust [55] and TrustLLM [40] further structure privacy evaluation along two complementary dimensions, privacy awareness and privacy leakage, providing a more systematic assessment framework.
A natural extension of these benchmarks is to combine privacy awareness and privacy leakage into a unified evaluation framework that jointly covers personal privacy, trade secrets, and state secrets. Such a naive fusion can be viewed as a comprehensive benchmark that evaluates both whether models recognize privacy-sensitive inputs and whether they disclose sensitive information in their outputs. However, this straightforward combination introduces several methodological limitations. First, it exhibits an inherent bias toward conservative models: models that frequently refuse to answer tend to achieve higher scores on the privacy leakage dimension, even at the cost of substantially degrading normal interaction quality. As a result, the evaluation framework implicitly favors over-refusal behaviors rather than balanced privacy protection mechanisms. Second, existing fusion-style benchmarks typically organize privacy leakage from task-oriented perspectives rather than from the standpoint of underlying model capabilities. For example, MultiTrust [55] categorizes privacy leakage into vision-based PII queries, visual-side privacy leakage, and conversational privacy leakage. Such task-driven categorizations may lead to overlapping evaluation dimensions and incomplete coverage, as they do not explicitly disentangle the distinct model abilities involved in privacy violations.
These limitations suggest the necessity of a more principled benchmark design that corrects evaluation bias and aligns privacy leakage dimensions with fundamental model capabilities, especially for large vision-language models.
@@ III-A Definition of Privacy
Past category-level privacy benchmarks lack a formalized privacy definition. Some category-level privacy benchmarks [29, 43, 55] focus on the security of personal identifiable information, which is only a subset of privacy. MLLMGUARD [17] expands the scope of privacy to personal privacy, trade and state secrets, but it still lacks a clear definition of privacy.
@@ IV Multi-PA Benchmark
Figure 3: VQA Generation Process. For security reasons, we obscure the private parts in images. We build Image Database and Attribute Database by collecting images and designing attributes for each privacy category. For each task, we create a variety of question templates which will be randomly selected to generate samples. Each VQA sample is the combination of an image from Image Database and a question from Question Templates. For Privacy Question Detection and Privacy InfoFlow Assessment, context of each sample is respectively from corresponding question in Privacy Leakage and sample in CONFAIDE [29].
@@ IV-A Dataset Overview
The design of Multi-PA is explicitly guided by the research hypotheses introduced in Section I. Specifically, the benchmark is constructed to enable the joint analysis of privacy awareness and privacy leakage (P1), to disentangle privacy leakage arising from distinct model capabilities (P2), and to investigate potential biases across different privacy types (P3). As shown in Table 2, we evaluate the privacy security of LVLMs from Privacy Awareness and Privacy Leakage.
@@
Privacy Question Detection. Privacy Question Detection assesses the model’s capacity to discern the privacy risks of input queries, which is specifically designed to test P1: whether a model’s awareness of the privacy sensitivity of a query is a necessary condition for mitigating privacy leakage. Notably, an image with privacy-related visual cues may be paired with a benign question, while a privacy-unrelated image may be associated with a sensitive question (e.g., querying for a politician’s home address given their publicly available photo is undeniably inappropriate). Therefore, the model’s assessment on the privacy sensitive of input queries may be context-dependent. To comprehensively evaluate a model’s awareness on the privacy sensitivity of input questions, we construct a dataset comprising image-question pairs, where images may contain private visual cues or not, and questions may be either privacy-related or privacy-unrelated. The model is tasked with determining the privacy relevance of each query.
@@
Privacy Leakage. Privacy leakage occurs when a model, during its interaction with users, either intentionally or unintentionally reveals privacy-related data. We categorize privacy leakage into two paradigms: assistant-like privacy leakage and database-like privacy leakage. Assistant-like leakage arises from insecure requests, where the model is compelled to disclose private data based on the provided input images. Database-like leakage on the other hand, occurs when private information, potentially present in the model’s training data, is memorized later exposed during interactions. Furthermore, we divide assistant-like leakage into “Perception” Leakage and “Reasoning” Leakage based on the capacities the model uses in the leakage process, and using “Memory” Leakage to evaluate the model’s database-like leakage. This categorization is capability-oriented rather than task-oriented, as each leakage type corresponds to a distinct underlying model capability, enabling explicit testing of P2. Additionally, we generate a variety of requests that do not involve private information, recorded as “Insensitive Questions”, which are used to assess the potential negative impact of privacy protection mechanisms on the model’s responsiveness to benign inquiries.
@@
Reasoning Leakage. “Reasoning” Leakage evaluates the risk of LVLMs inferring privacy based on visual cues within input images. Recent researches have demonstrated language models can infer privacy-sensitive data from indirect clues [39, 43]. Motivated by these findings, we consider the ability of models to infer privacy attributes as a crucial aspect of model’s privacy risks. To prevent models from facilitating the illicit prediction of private data, a privacy-preserving model should avoid generalizing its reasoning abilities to inferring privacy attributes. Therefore, for each privacy category, we curate a series of inferable privacy attributes for each privacy category. These attributes can be indirectly deduced from cues present in input images. For example, in an image containing a company’s revenue data, the company’s financial status might be considered as an inferable attribute.
Memory Leakage. “Memory” Leakage evaluates the risk of LVLMs revealing privacy retained from its training process. Researches have shown that models can memorize training data and unintentionally disclose it during interactions [7, 9, 18, 53, 39]. Privacy with long-term sensitivity, like PII of public figures, may be unintentionally exposed on social platforms and subsequently incorporated into model training datasets. The privacy leakage of such information may lead to detrimental consequences. To assess the risk of a model leaking private information from its training dataset, we define weakly associated privacy attributes for each privacy category. Weakly associated privacy attributes refer to information that cannot be directly or indirectly derived from the input data. However, models may still output these attributes based on private information it has memorized from its training dataset. For example, given images of some businessmen, we ask models about their home addresses. Such private data may have appeared in journalists’ reports. When these reports are used for model training, models might memorize such private data and provide correct answer during inference.